OpenStudy (anonymous):

need help,Please! square{17} x cube root 17^2

OpenStudy (anonymous):

square means power of 2 cube root means power of 1/3.. can you solve now?

OpenStudy (anonymous):

\[\sqrt{17}\times \sqrt[3]{17^2}\]

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

so it is \[17^{1/2}\times17^{2/3}\]

OpenStudy (anonymous):

the exponents get added together if the base is the same number

OpenStudy (anonymous):

so what is 1/2 + 2/3?

OpenStudy (anonymous):

5/6

OpenStudy (anonymous):

than the answer is \[17^{5/6}\]

OpenStudy (anonymous):

or \[\sqrt[6]{17^{5}}\]

OpenStudy (anonymous):

ok, I get it now,TY

OpenStudy (anonymous):

your welcome